A bag is filled with green and blue marbles. There are 48 marbles in the bag. If there are 38 more green marbles than blue marbles, find the number of green marbles and the number of blue marbles in the bag.
Find the solution by setting up an equation. You have two choices for how you will pick your variable. For each of the following variable choices, write an equation (just set up the equation, do not solve it yet) that describes the relationship between the number of green marbles, the number of blue marbles, and the total number of marbles in the bag:
If x = number of green marbles in the bag, my equation is = 48
If x = number of blue marbles in the bag, my equation is = 48
Now, solve one of your equations to determine the number of green and number of blue marbles:
There are green marbles and blue marbles.

Answer by stanbon(75887)

A bag is filled with green and blue marbles. There are 48 marbles in the bag. If there are 38 more green marbles than blue marbles, find the number of green marbles and the number of blue marbles in the bag.
------
Equations:
g + b = 48 marbles
g = b + 38
------
Substitute for "g" and solve for "b"::
b+38 + b = 48
2b = 10
b = 5 (# of blue marbles)
g = b+38 = 43 (# of green marbles)
-------------
Cheers,
Stan H.
----------